# Analysis of ion chain sympathetic cooling and gate dynamics

### v1

@@ III.3 Cooling optimization
For a cooling Rabi frequency of 640 kHz, we find that an optimal cooling duty cycle consists of one gate and a $\approx$75cooling duty cycle. This results in a total circuit fidelity of 0.73 and a mean gate fidelity of 0.9993. As seen in Fig. 4, as cooling duration is increased, average gate fidelity increases quickly when the duty cycle is short enough to address motional errors before slowly dropping off due to T2 times. In addition, the fidelity decreases as the number of gates per cycle is increased, as increasing the number of gates per cycle leads to more time for motional error to accrue. In all cases, we observed that optimal fidelity was achieved when cooling after every gate run in the circuit, as seen in Figure 4. This corresponds to the fact that $F\propto\frac{1}{\sqrt{1+n^{2}}}$ Cetina et al. (2022a). Thus, for small $n$ ($n\approx 0$), $dF/dn\approx 0$, meaning that by running only one gate at a time, the errors due to heating do not have time to manifest.
@@
It is important to also note that in this analysis we do not account for photon recoil heating of the radial modes. It was shown in Cetina et al. (2022a) that these modes could also be cooled using the same methods as the COM mode, and in a similar amount of time. Therefore, we can assume that cooling the radial modes would add overhead to the cooling process, although the exact optimal scheme for cooling all modes is outside the scope of this paper. In Appendix C, we show that changing the cooling time by a factor of two to account for this overhead shifts the results in Fig. 4(b) towards higher performance at shorter cooling times, but do not change the overall behavior in this parameter regime.
@@ IV Outlook
As discussed in Cetina et al. (2022b), the operation of long ion chains for quantum computation is primarily limited by motional heating along the axis of the chain. A solution is proposed to used sympathetic coolants, and in this study we expand that idea and optimize coolant placement and number to maximize circuit fidelity given certain experimental parameters. We have formulated a first-order approximation of cooling dynamics to analytically show that coolant placement at the center of the chain is optimal, placing an upper bound on the cooling limit of the COM mode of an ion chain. Our theoretical findings show that improvements in heating rate allow for a much more significant decrease in cooling limit, and therefore optimal circuit fidelity. Unsurprisingly, we find that the number of coolants needed for a chain must strike a balance between increasing axial heating due to a decrease in COM mode frequency and increasing cooling rate due to stronger coupling to the COM mode.
@@
This study examines the influence of system improvements on the performance of various cooling schemes; in particular, this study examines how improvements in cooling power and cooling rate can lead to improvements in circuit fidelity. However, further study is required regarding improvements of other aspects of the device, such as decreasing heating rates Hite et al. (2012); Daniilidis et al. (2014) or using open-loop control methods such as dynamical decoupling to achieve a longer T2 time Viola and Lloyd (1998); Biercuk et al. (2009).
@@ V Acknowledgments
APPENDIX A TRAP FREQUENCY AND COOLANT RATIOS
Figure 5: Cooling performance and trap frequency. Cooling limit with respect to coolant proportion and COM frequency. The increase in cooling limit in the vertical direction corresponds with an increase in heating rate due to a lower COM frequency, while the decrease in cooling limit in the rightward direction corresponds to a larger number of coolants in the chain. The cooling limit plateaus as the chain fills with coolants.
For the purposes of this study, we assume a fixed electric potential for ion trapping. However, in this section, we briefly examine how changing aspects of the trap potential (in particular, the COM mode frequency) impacts cooling performance.
In addition, rather than adding coolants to the chain for a fixed number of qubits, we examine the ratio of coolants to total ions in the chain. In so doing, we attempt to examine the effect of increasing the cooling rate of the mode by maximizing the total mode participation over all coolants (i.e., maximizing $\sum_{k\in P}|v_{0k}|^{2}$ in equation 3).
As such, we test 21-ion chains with differing COM frequencies as we add coolants starting at the center of the chain, subsequently assessing their cooling performance. Figure 5 shows the corresponding cooling limit of a chain for a fixed COM frequency and coolant filling fraction. The theoretical prediction for our cooling limit holds well here, with the cooling limit decreasing as the COM frequency increases. Thus, with ion motion more restricted by the trap potential, heating is suppressed, allowing for easier cooling. In addition, the cooling performance plateaus as the coolant filling fraction approaches 1. This is indicative of the trends observed in Figure 3, in which cooling performance plateaus and eventually decreases as chain length increases.
APPENDIX B COMPUTATIONAL VERIFICATION OF COOLANT PLACEMENT
Figure 6: Cooling performance and coolant positions. Simulated COM mode cooling rates for all possible configurations of two coolants in a 15-ion chain. Optimal placement is found with ions placed at positions (0, 1) or (-1, 0), in good agreement with our theoretical observation that optimal cooling performance occurs with ions placed at the center of the chain.
Figure 7: Cooling dynamics with cooling limit. Cooling dynamics with illustrated cooling limit. The analytical cooling limit shows good agreement with computational results as well.
Figure 8: Cooling parameter analysis with radial cooling. Cooling results when accounting for radial cooling. (a) Mean gate fidelity with respect to gate time and cooling time per cycle. (b) Total circuit fidelity with respect to gate time and cooling time per cycle. (c) Optimal circuit fidelity with respect to number of gates per cycle. The optimal configuration still occurs when only one gate is run per cycle, while the optimal duty cycle is slightly smaller than when not accounting for photon recoil.
We also use computational methods to verify the accuracy of our analytical observations by examining the correspondence between the perturbative definitions of the cooling rate and cooling limit as defined in section II. Figure 7 shows the accuracy of the cooling limit. Our analytical results show good agreement to the simulated dynamics, accurately estimating the limit of a particular cooling scheme given its geometry. In addition, we also found that the perturbative approximation of the cooling rate only had a relative error of $10^{-4}$ for $\gamma<10^{4}$.
In section II, we perturbatively show that the COM mode is cooled most effectively when maximizing mode participation factors. As such, placing coolant ions at the center of the chain maximally couples their motion to the COM mode, thus yielding optimal cooling performance.
We verify this computationally by simulating every possible configuration of two coolant ions in a 15-ion chain, seen in figure 6. Cooling performance was maximized with centrally-placed coolants (i.e. at positions (-1, 0) or (0, 1)), yielding a cooling limit of 29 quanta. This is in good agreement with our theoretical observations, as configurations that perform best (i.e. minimize the cooling limit) are centrally placed, while configurations that perform worse place coolant ions towards the edge of the chain.
APPENDIX C PHOTON RECOIL AND COOLING
When accounting for photon recoil, we add a wait time to account for radial cooling time. During this wait, we neither run gates nor cool. We assume that an optimal cooling scheme could be constructed empirically to ensure all modes are close to the cooling limit at the end of the cooling time, such as alternating which mode is cooled over time.
Results from these modified simulations can be seen in figure 8. Even accounting for photon recoil, our salient observations did not change; optimal fidelity was still achieved when cooling as often as possible. Although the amount of time required for optimal performance changed, it didn’t change significantly: a Rabi frequency of 180 kHz yielded optimal results at a duty cycle of 84.82% (1258 $\mu$s of cooling time per cycle), a 275 kHz Rabi frequency yielded an optimal duty cycle of 70.84% (547 $\mu$s of cooling time per cycle), while at 640 kHz, a duty cycle of 65.54% was found to be optimal (428 $\mu$s of cooling time per cycle).
Each of these optimal duty cycles was found to be slightly lower than the corresponding results that do not account for radial cooling time; this is most likely due to the fact that extra radial cooling leads to longer circuits, which causes dephasing errors to become more prevalent in the circuit error budget. As such, reducing the cooling time slightly can trade off heating performance for a shorter circuit, and thus smaller dephasing errors. This can likewise be seen in Figure 8(b), with increased performance for configurations with smaller cooling duty cycles.

### v2

@@ III.3 Cooling optimization
For a cooling Rabi frequency of 640 kHz, we find that an optimal cooling duty cycle consists of one gate and a $\approx$75cooling duty cycle. This results in a total circuit fidelity of 0.73 and a mean gate fidelity of 0.9993. As seen in Fig. 5, as cooling duration is increased, average gate fidelity increases quickly when the duty cycle is short enough to address motional errors before slowly dropping off due to T2 times. In addition, the fidelity decreases as the number of gates per cycle is increased, as increasing the number of gates per cycle leads to more time for motional error to accrue. In all cases, we observed that optimal fidelity was achieved when cooling after every gate run in the circuit, as seen in Fig. 5. This corresponds to the fact that $F\propto\frac{1}{\sqrt{1+n^{2}}}$ [3]. Thus, for small $n$ ($n\approx 0$), $dF/dn\approx 0$, meaning that by running only one gate at a time, the errors due to heating do not have time to manifest.
@@
It is important to also note that in this analysis we do not account for photon recoil heating of the radial modes. It was shown in [3] that these modes could also be cooled using the same methods as the COM mode, and in a similar amount of time. Therefore, we can assume that cooling the radial modes would add overhead to the cooling process, although the exact optimal scheme for cooling all modes is outside the scope of this paper. During this added cooling overhead, we neither run gates nor cool. We assume that an optimal cooling scheme could be constructed empirically to ensure all modes are close to the cooling limit at the end of the cooling time, such as alternating which mode is cooled over time.
When accounting for photon recoil, optimal fidelity was still achieved when cooling as often as possible. The amount of time required for optimal performance changed slightly: a Rabi frequency of 180 kHz yielded optimal results at a duty cycle of 84.82% (1258 $\mu$s of cooling time per cycle), a 275 kHz Rabi frequency yielded an optimal duty cycle of 70.84% (547 $\mu$s of cooling time per cycle), while at 640 kHz, a duty cycle of 65.54% was found to be optimal (428 $\mu$s of cooling time per cycle). Each of these optimal duty cycles was found to be slightly lower than the corresponding results that do not account for radial cooling time. This descrease is most likely due to the fact that extra radial cooling leads to longer circuits, which causes dephasing errors to become more prevalent in the circuit error budget. Reducing the cooling time slightly can trade off heating performance for a shorter circuit, and thus smaller dephasing errors. This leads to generally increased performance for configurations with smaller cooling duty cycles.
@@ IV Outlook
As discussed in [10], the operation of long ion chains for quantum computation is primarily limited by motional heating along the axis of the chain. A solution is proposed to used sympathetic coolants, and in this study we expand that idea and optimize coolant placement and number to maximize circuit fidelity given certain experimental parameters. We have formulated a first-order approximation of cooling dynamics to analytically show that coolant placement at the center of the chain is optimal, placing an upper bound on the cooling limit of the COM mode of an ion chain. Our theoretical findings show that improvements in heating rate allow for a much more significant decrease in cooling limit, and therefore optimal circuit fidelity. Unsurprisingly, we find that the number of coolants needed for a chain must strike a balance between increasing axial heating due to a decrease in COM mode frequency and increasing cooling rate due to stronger coupling to the COM mode.
@@
This study examines the influence of system improvements on the performance of various cooling schemes; in particular, this study examines how improvements in cooling power and cooling rate can lead to improvements in circuit fidelity. However, further study is required regarding improvements of other aspects of the device, such as decreasing heating rates [21, 22] or using open-loop control methods such as dynamical decoupling to achieve a longer T2 time [23, 24].